• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 00:38:29 on October 31,1998

    Entry number 4225

    keyword=target NMR

    11:57 NMR Measurements

    UP: -14.08mV (fitted)
    DN: -14.13mV (fitted)
    --------->38.09% EPR
    --------->36.67% H2O

    Note: MCC has told us that we will not have beam for ~1/2 hr so took the opportunity
    to make another measurement. This measurement confirms my earlier
    suspicion that the current to the coils is not completely settled after
    field rotation.
    Position of UP Resonance :
    ---> 27.65 G (immediately after rotation)
    ---> 27.80 G (~2hrs later)

    ----------------------------------------------------

    23:51 moved target into NMR position
    23:54 NMR/RTD screen shot.
    23:58 target back to beam position (39246)